# FastVLM: Efficient Vision Encoding for Vision Language Models

### v1

@@ 1-2B Model Comparison
R41 Cambrian-1 [72] ViT-L/14 L-3 2.5+7 336 576 304 5085 64.6 80.4 71.7 - - - - 77.8 42.7 74.7
Table 6: VLM evaluations and comparison with recent methods. The models are grouped based on total number of visual tokens. “-” indicates that performance was not reported in the respective paper. For the dataset column, “-” indicates that the dataset size for pretraining (“PT”) or instruction tuning (“IT”) is not explicitly mentioned in the respective paper. For methods that have more than 2 stages of training, we report the total samples used for all the pretraining stages as part of “PT”. “TTFT” means time to first token (the sum of the vision encoder latency and the LLM prefilling time), we report latency only for models that are publicly available and in a format favorable to MLX [27] “Vic.” refers to Vicuna [90], “Qw.2” refers to Qwen2 [79] and “Qw.” refers to Qwen [3]. “L-2” refers to LLaMA-2. “L-3” refers to LLaMA-3. “ML.” refers to MobileLLaMA [17, 18]. “DS.” refers to DeepSeek LLM [20]. For input resolution and visual tokens, we report the highest supported resolution by the respective models as some models like LLaVA-OneVision [40] and MM1 [60] use dynamic input resolution. †- performance numbers reported from [72]. For MiniGeminiHD [44], the dataset sizes is inferred from the pretrain and instruction tuning JSON files. For VLMs that use multiple vision encoders, the size of each encoder is listed independently, for TTFT, the latency from each encoder is summed up.
@@ 4 Experiments
In this section, we present our training setup and results.
@@
In Sec. 4, we present results with different LLM decoders, primarily with Qwen2-0.5B/1.5B/7B model family [79] (chat variant) and Vicuna-7B model [90]. We report results in two training setups, the first one is the 2-Stage setup introduced in LLaVA-1.5. We additionally scale the dataset used in Stage 2 from 665k samples to 1.1 million samples, which is a subset of the instruction tuning dataset used in InternVL [14] (more details in Sec. D). For the second training setup, we follow the current trend in literature [38, 60] of training the VLMs in 3 stages, i.e. Stage 1 for training the connector, Stage 1.5 for resolution scaling and Stage 2 for visual instruction tuning. In Stage 1.5, we use the densely captioned CC3M [65] and CC12M [10] datasets [38]. For the final stage we use 1.1 million instruction tuning dataset. In this setup, the input image resolution is set to the backbone pretraining resolution for Stage 1 and adjusted to the target resolution for the following two stages. In both setups, the vision encoder and LLM are frozen only in stage 1, while all modules are finetuned in the remaining stages.
@@
For ablations presented in Sec. 3, we report GQA, TextVQA, POPE, DocVQA and SeedBench. GQA and SeedBench are general knowledge benchmarks, DocVQA and TextVQA represent text-rich evaluations and POPE is a hallucination benchmark. Together these benchmarks provide diversity and are quick to evaluate for ablations. Most importantly, they exhibit lower variance to different initializations and under probabilistic decoding setting. We report the variance for all the evals for different initialization in Sec. D.3. The standard deviation across the 5 selected metrics is less than 0.5. We call the average of these 5 benchmarks Avg-5, and use it as a reliable signal for our analysis. Our empirical estimate of the standard deviation for Avg-5 is $\simeq$0.1.
@@ 4.1 Comparison with state-of-the-art
Hierarchical Backbones. When we compare FastVLM (R18) with ConvLLaVA [25] (R16), with the same LLM and similar training data size, our model obtains +8.4% better performance on TextVQA and +12.5% better performance on DocVQA while being 22% faster. The gap widens at higher resolution, where FastVLM (R26 and R27) achieves superior performance on wide range of benchmarks while being 2$\times$ faster than ConvLLaVA (R24), with the same LLM decoder.
@@
Dataset Scaling. When designing a new architecture, it is crucial to consider how effectively the model scales with training data. We demonstrate the performance of FastVLM when scaling the pretraining and instruction tuning datasets. By increasing the instruction tuning dataset from 0.6M to 1.1M samples (R18), FastVLM outperforms prior works like MobileVLMv2 [18] (R15), which was trained on larger instruction tuning and pretraining datasets. Additionally, when scaling the pretraining dataset by incorporating an intermediate pretraining stage for resolution scaling with 15M samples, FastVLM (R19) matches or surpasses MM1 [60] (R36) across a wide range of benchmarks, including MMMU, MMVet, SQA, POPE, and SeedBench. Remarkably, FastVLM achieves this performance while generating 5$\times$ fewer visual tokens. With an input resolution of 1024$\times$1024 and a larger instruction tuning dataset of size 11.9M, FastVLM (R39) outperforms MM1 (R36) and LLaVA-NeXT (R37) across various benchmarks. Even on text-rich evaluations, like TextVQA and DocVQA, which are sensitive to input resolution and number of visual tokens, our model achieves better performance with 2.8$\times$ and 11.3$\times$ less visual tokens than MM1 and LLaVA-NeXT respectively. We provide details of the dataset splits in Sec. D.
@@
Multiple Vision Encoders. Recently, MiniGemini [44] and Cambrian-1 [72] introduced models that rely on multiple vision encoders. In Tab. 6, we compare FastVLM (R38), which uses a single vision encoder with methods that use multiple encoders and trained on similarly scaled visual instruction tuning dataset. In Cambrian-1 [72] (R41), vision encoding contributes 3.2$\times$ more than LLM prefilling to the total time-to-first-token of approximately 5 seconds (detailed breakdown is provided in Tab. 9). FastVLM (R38) outperforms Cambrian-1 (R41) when trained on a similar visual instruction tuning dataset, while being 7.9$\times$ faster. By scaling the instruction tuning dataset to 11.9M, FastVLM (R39) achieves superior performance over Cambrian-1 (R41) with 2.3$\times$ fewer visual tokens, even on text-rich evaluations (see Tab. 10) that are sensitive to the number of visual tokens.
@@
Effect of Decoder. VLM performance also depends on the quality of LLM, as demonstrated in prior studies, like [39]. By switching from Vicuna-7B (R19, R27) to Qwen2 [79, 71] models (R20, R28), we see a good improvement in performance across all the benchmarks. The improvements are significant on MMVet, LLaVA-in-the-wild and MMMU benchmarks. With Qwen2-0.5B as the LLM decoder, FastVLM (R3) matches the performance of LLaVA-OneVision [40] (R2) on key benchmarks such as SeedBench, MMMU, and MMVet, while being 85$\times$ faster and trained on 2.9$\times$ fewer instruction tuning samples. This result underscores the quality of our vision encoder, as both models use the same LLM decoder, while FastViTHD is 3.4$\times$ smaller compared to SigLIP-SO400M [87].
@@ 5 Conclusion
In this work, we introduced FastVLM, which leverages FastViTHD image encoders designed for enhanced resolution scaling while maintaining efficiency. By strategically trading off the costly self-attention in ViTs with a purpose-built hybrid architecture, FastViTHD processes high-resolution images efficiently, and outputs a substantially reduced number of visual tokens. The design of FastVLM enables competitive performance with prior works across a wide range of VLM benchmarks, while improving efficiency in both time-to-first-token and the number of parameters in the vision backbone. Rigorous benchmarking on an M1 MacBook Pro demonstrates that FastVLM achieves a state-of-the-art resolution-latency-accuracy trade-off compared to existing works.
@@ A Training Setup
To showcase our model’s performance in the presence of additional dataset, we scale both pretraining and instruction tuning datasets in Sec. 4. For results presented in R13, R17, R18, R25, R26 in Tab. 6, we still perform 2-stage training described in Tab. 7, for R18 and R26, we use instruction tuning dataset of size 1.1 million samples in Stage-2. For results presented in R3, R4, R6, R7, R10, R11, R14, R19, R20, R21, R27, R28, R38 and R39, we scale-up both instruction tuning dataset and pretraining dataset. We also introduce and additional stage of pretraining with the scaled-up dataset as described in Tab. 8. Details of 1.1 million, 6.5 million and 11.9 million instruction tuning dataset is presented in Sec. D.
@@
Stage-1 Stage-1.5 Stage-2
Data LLaVA-1.5 558K Recap-CC3M + 1.1M / 6.5M / 11.9M
Recap-CC12M [38]
Learning Rate 1e-3 2e-5 2e-5
Batch size 256 128 128
LR. schedule cosine decay cosine decay cosine decay
LR. warmup ratio 0.03 0.03 0.03
Optimizer AdamW AdamW AdamW
Trainable Projector Full Full
modules Model Model
Table 8: 3-Stage training setup used for results with scaled-up data in Tab. 6.
@@ 1-2B Model Comparison
Table 9: Breakdown of prefilling latencies for recent methods. The models are grouped based on total number of visual tokens. For models that were difficult to export or unavailable, we mark them as ’-’ in the table. “Vic.” refers to Vicuna [90], “Qw.2” refers to Qwen2 [79] and “Qw.” refers to Qwen [3]. “L-2” refers to LLaMA-2. “L-3” refers to LLaMA-3. “ML.” refers to MobileLLaMA [17, 18]. “DS.” refers to DeepSeek LLM [20]. For input resolution and visual tokens, we report the highest supported resolution by the respective models as some models like LLaVA-OneVision [40] and MM1 [60] use dynamic input resolution. For VLMs that use multiple vision encoders, the size of each encoder is listed independently, for TTFT, the latency from each encoder is summed up.
@@ C Additional Results
R1 MM1 [60] ViT-H - 3000+1.5 1344 720 632 - 72.6 62.6 72.8 76.8 45.5
R2 LLaVA-NeXT ViT-L/14 L-3 - 672 2880 304 20347 69.5 49.0 64.6 72.6 -
R3 Cambrian-1 [72] ViT-L/14 L-3 2.5+7 336 576 304 5085 73.3 62.4 71.7 77.8 -
@@
R4 FastVLM (Ours) FastViTHD Vic. 0.5+0.6 768 144 125 387 17.1 30.0 62.9 32.9 28.7
R5 FastVLM (Ours) FastViTHD Vic. 0.5+1.1 768 144 125 387 59.1 38.4 67.5 57.3 29.7
R6 FastVLM (Ours) FastViTHD Vic. 15+1.1 768 144 125 387 65.4 45.3 69.4 65.5 32.0
R7 FastVLM (Ours) FastViTHD Qw.2 15+1.1 768 144 125 446 69.3 45.9 69.5 66.9 34.3
R8 FastVLM (Ours) FastViTHD Qw.2 15+11.9 768 144 125 446 74.2 59.0 72.8 72.0 44.3
R9 FastVLM (Ours) FastViTHD Vic. 0.5+0.6 1024 256 125 577 19.2 29.3 64.4 35.6 28.9
R10 FastVLM (Ours) FastViTHD Vic. 0.5+1.1 1024 256 125 577 61.0 38.3 67.4 62.8 32.0
R11 FastVLM (Ours) FastViTHD Vic. 15+1.1 1024 256 125 577 66.9 47.1 70.6 72.4 34.7
R12 FastVLM (Ours) FastViTHD Qw.2 15+1.1 1024 256 125 641 71.0 49.7 72.1 73.3 37.5
R13 FastVLM (Ours) FastViTHD Qw.2 15+6.5 1024 256 125 641 76.6 52.9 73.1 78.7 44.2
R14 FastVLM (Ours) FastViTHD Qw.2 15+11.9 1024 256 125 641 77.0 63.3 74.8 78.9 49.7
@@
Table 10: Comparison with recent methods on text-rich benchmarks. The models are grouped based on total number of visual tokens. “-” indicates that performance was not reported in the respective paper. For the dataset column, “-” indicates that the dataset size for pretraining (“PT”) or instruction tuning (“IT”) is not explicitly mentioned in the respective paper. For methods that have more than 2 stages of training, we report the total samples used for all the pretraining stages as part of “PT”. “TTFT” means time to first token (the sum of the vision encoder latency and the LLM prefilling time), we report latency only for models that are publicly available and in a format favorable to MLX [27] “Vic.” refers to Vicuna [90], “Qw.2” refers to Qwen2 [79]. “L-3” refers to LLaMA-3. * - For input resolution and visual tokens, we report the highest supported resolution by the respective models as some models like MM1 [60] use dynamic input resolution. †- performance numbers reported from [72]. For VLMs that use multiple vision encoders, the size of each encoder is listed independently, for TTFT, the latency from each encoder is summed up.

### v2

@@ 1-2B Model Comparison
R44 Cambrian-1 [78] ViT-L/14 L-3 2.5+7 336 576 304 5085 64.6 80.4 71.7 - - - - 77.8 42.7 74.7
Table 6: VLM evaluations and comparison with recent methods. The models are grouped based on total number of visual tokens. “-” indicates that performance was not reported in the respective paper. For the dataset column, “-” indicates that the dataset size for pretraining (“PT”) or instruction tuning (“IT”) is not explicitly mentioned in the respective paper. For methods that have more than 2 stages of training, we report the total samples used for all the pretraining stages as part of “PT”. “TTFT” means time to first token (the sum of the vision encoder latency and the LLM prefilling time), we report latency only for models that are publicly available and in a format favorable to MLX [31] “Vic.” refers to Vicuna [98], “Qw.2” refers to Qwen2 [86] and “Qw.” refers to Qwen [3]. “L-2” refers to LLaMA-2. “L-3” refers to LLaMA-3. “ML.” refers to MobileLLaMA [18, 19]. “DS.” refers to DeepSeek LLM [21]. For input resolution and visual tokens, we report the highest supported resolution by the respective models as some models like LLaVA-OneVision [45] and MM1 [66] use dynamic input resolution. FastVLM models using dynamic resolution employs a simple 2$\times$2 grid, with tile size set to 1024. †- performance numbers reported from [78]. For VLMs that use multiple vision encoders, the size of each encoder is listed independently, for TTFT, the latency from each encoder is summed up.
@@ 4 Experiments
In Sec. 4, we present results with different LLM decoders, primarily with Qwen2-0.5B/1.5B/7B model family [86] (chat variant) and Vicuna-7B model [98]. We report results in two training setups, the first one is the 2-Stage setup introduced in LLaVA-1.5. For the second training setup, we follow the current trend in literature [43, 66] of training the VLMs in 3 stages, i.e. Stage 1 for training the connector, Stage 1.5 for resolution scaling and Stage 2 for visual instruction tuning. Information on datasets used in these stages can be found in Sec.D. In this setup, the input image resolution is set to the backbone pretraining resolution for Stage 1 and adjusted to the target resolution for the following two stages. In both setups, the vision encoder and LLM are frozen only in stage 1, while all modules are finetuned in the remaining stages. For the best setup, we further finetune the model from Stage 2 with high quality instruction tuning dataset with chain-of-thought reasoning from [30] and call this Stage 3. Details of this setup is elaborated in Sec. A and Sec. D. We publicly release R4, R12, and R41 checkpoints from Stage 2 training, as well as the R5, R13, and R42 checkpoints from Stage 3 training, as part of our open-sourced codebase.
@@
For ablations presented in Sec. 3, we report GQA, TextVQA, POPE, DocVQA and SeedBench. GQA and SeedBench are general knowledge benchmarks, DocVQA and TextVQA represent text-rich evaluations and POPE is a hallucination benchmark. Together these benchmarks provide diversity and are quick to evaluate for ablations. Most importantly, they exhibit lower variance to different initializations and under probabilistic decoding setting. We report the variance for all the evals for different initialization in Sec. D.3. The standard deviation across the 5 selected metrics is less than 0.5. We call the average of these 5 benchmarks Avg-5, and use it as a reliable signal for our analysis. The empirical standard deviation estimate for Avg-5 is 0.1.
@@ 4.1 Comparison with state-of-the-art
Hierarchical Backbones. When we compare FastVLM (R20) with ConvLLaVA [28] (R18), with the same LLM and similar training data size, our model obtains +8.4% better performance on TextVQA and +12.5% better performance on DocVQA while being 22% faster. The gap widens at higher resolution, where FastVLM (R28 and R29) achieves superior performance on wide range of benchmarks while being 2$\times$ faster than ConvLLaVA (R26), with the same LLM decoder.
@@
Dataset Scaling. When scaling the pretraining dataset by incorporating an intermediate pretraining stage for resolution scaling with 15M samples, FastVLM (R21) matches or surpasses MM1 [66] (R38) across a wide range of benchmarks. Remarkably, FastVLM achieves this performance while generating 5$\times$ fewer visual tokens. With an input resolution of 1024$\times$1024 and a larger instruction tuning dataset of size 12.5M, FastVLM (R41) outperforms MM1 (R38) and LLaVA-NeXT (R39) across various benchmarks, including text-rich evaluations, like TextVQA and DocVQA, which are sensitive to input resolution and number of visual tokens. The gap widens when further scale up input resolution using AnyRes, more details in Sec. C.1. We provide details of the dataset splits in Sec. D.
@@
Multiple Vision Encoders. Recently, MiniGemini [49] and Cambrian-1 [78] introduced models that rely on multiple vision encoders. In Tab. 6, we compare FastVLM (R40), which uses a single vision encoder with methods that use multiple encoders and trained on similarly scaled visual instruction tuning dataset. In Cambrian-1 [78] (R44), vision encoding contributes 3.2$\times$ more than LLM prefilling to the total time-to-first-token of approximately 5 seconds (detailed breakdown is provided in Tab. 10). FastVLM (R40) outperforms Cambrian-1 (R44) when trained on a similar visual instruction tuning dataset, while being 7.9$\times$ faster. By scaling the instruction tuning dataset to 12.5M, FastVLM (R41) achieves superior performance over Cambrian-1 (R44) with 2.3$\times$ fewer visual tokens, even on text-rich evaluations (see Tab. 11) that are sensitive to the number of visual tokens.
@@
Effect of Decoder. VLM performance also depends on the quality of LLM, as demonstrated in prior studies, like [44]. By switching from Vicuna-7B (R21, R29) to Qwen2 [86, 77] models (R22, R30), we see improvement in performance across all the benchmarks. The improvements are significant on MMVet, LLaVA-in-the-wild and MMMU benchmarks. With Qwen2-0.5B as the LLM decoder, FastVLM (R4) outperforms LLaVA-OneVision [45] (R2) while being 85$\times$ faster. This result underscores the quality of our vision encoder, as both models use the same LLM decoder, while FastViTHD is 3.4$\times$ smaller compared to SigLIP-SO400M [94].
Row Method Vision LLM #Visual Vis. Enc. ChartQA OCRBench TextVQA DocVQA InfoVQA MMMU AI2D SQA Seed
Ann. Encoder Tokens Size(M)$\downarrow$ BenchI
0.5B Model Comparison
R1 SmolVLM2-0.5B [62] ViT-B SmolLM2 1088 93 62.8 61.0 60.2 70.5 25.5 33.7 59.2 80.0 62.2
R2 FastVLM-0.6B (Ours) FastViTHD Qw.2 256 125 71.4 55.8 65.8 79.1 43.3 33.6 66.0 82.0 69.3
1-3B Model Comparison
R3 SmolVLM2-2.2B [62] ViT-SO400M SmolLM2 2106 430 68.7 72.9 73.0 80.0 37.8 42.0 70.0 89.6 71.3
R4 MolmoE A1B-7B [22] ViT-SO400M OLMoE. 1728 430 48.0 54.7 61.5 77.7 - 33.9 71.0 87.5 67.9
R5 MM1.5-1B [95] ViT-H - 1440 632 67.2 60.5 72.5 81.0 50.5 35.8 59.3 82.1 70.2
R6 FlorenceVL-3B [11] DaViT Phi-3 576 770 70.7 63.0 69.1 82.1 51.3 41.8 73.8 84.6 70.6
R7 FastVLM-1.7B (Ours) FastViTHD Qw.2 256 125 69.5 61.2 71.2 84.2 49.6 38.0 76.6 92.9 72.6
Table 7: Comparison with concurrent works. Models are ordered in the descending order of the number of visual tokens. The number of visual tokens listed for each model is based on the highest supported input resolution for the respective models. - For the number of visual tokens generated by SmolVLM2, we estimated it based on the model configs as there is no explicit mention of this information in the paper. The reported visual tokens correspond to the highest resolution supported by the respective model. - models are evaluated using VLMEvalKit [25]. The rest of the models are evaluated using lmms-eval [96] library. The FastVLM model used in this table is R5 and R13 from Tab. 6
Concurrent works. In Tab. 7, we compare FastVLM with concurrent works like SmolVLM2 [62] and FlorenceVL [11]. FastVLM (R7) performs competitively against SmolVLM2 (R3), particularly on text-rich evaluations such as ChartQA and TextVQA, while surpassing it on the DocVQA and InfoVQA benchmarks. FastVLM (R7) accomplishes this using 8.2$\times$ fewer visual tokens than SmolVLM2 (R3). Additionally, FastVLM (R7) outperforms FlorenceVL (R6) on benchmarks such as TextVQA and DocVQA, while using 2.3$\times$ fewer visual tokens and a 6.2$\times$ smaller vision encoder. On knowledge based evaluations like AI2D and ScienceQA, FastVLM (R7) outperforms all other competing models. At the smallest scale, FastVLM (R2) outperforms a similar sized SmolVLM2 model (R1) on a wide range of benchmarks while using 4.3$\times$ fewer visual tokens.
@@ 5 Conclusion
In this work, we introduced FastVLM, which leverages the FastViTHD vision backbone for efficient encoding of high-resolution inputs. FastViTHD has a hybrid architecture, is pretrained on reinforced image-text data, and outputs a substantially reduced number of visual tokens with minimal accuracy sacrifice. FastVLM has competitive performance with prior works across a wide range of VLM benchmarks while improving efficiency in both time-to-first-token and the number of parameters in the vision backbone. Rigorous benchmarking on an M1 MacBook Pro demonstrates that FastVLM achieves a state-of-the-art resolution-latency-accuracy trade-off compared to existing works.
Acknowledgement
We thank David Koski, Christopher Webb, Matt Biddulph, Nick Henderson, Megan Maher Welsh, Jamie Cheng, and Jerremy Holland for their valuable contributions to prototyping the demo app. We are also grateful to Angelos Katharopoulos and Awni Hannun for their insightful feedback and suggestions on benchmarking. Additionally, we thank Jen-Hao Rick Chang and Vishwanath Sindagi for their helpful feedback and recommendations.
@@ A Training Setup
To showcase our model’s performance in the presence of additional dataset, we scale both pretraining and instruction tuning datasets in Sec. 4. For results presented in R15, R19, R20, R27, R28 in Tab. 6, we still perform 2-stage training described in Tab. 8, for R20 and R28, we use instruction tuning dataset of size 1.1 million samples in Stage-2. For results presented in R3, R4, R7, R8, R11, R12, R16, R21, R22, R23, R29, R30, R40 and R41 we scale-up both instruction tuning dataset and pretraining dataset. We also introduce and additional stage of pretraining with the scaled-up dataset as described in Tab. 9. Details of 1.1 million, 6.5 million and 12.5 million instruction tuning dataset is presented in Sec.D. For results presented in R5, R13 and R42, we introduce additional finetuning on high quality instruction tuning dataset from [30] in Stage 3. We publicly release R4, R12, and R41 checkpoints from Stage 2 training, as well as the R5, R13, and R42 checkpoints from Stage 3 training, as part of our open-sourced codebase.
@@
Stage-1 Stage-1.5 Stage-2 Stage 3
Data LLaVA-1.5 558K Recap-CC3M + 1.1M / 6.5M 10.6M
Recap-CC12M [43] 12.5M
Learning Rate 1e-3 2e-5 2e-5 2e-5
Batch size 256 128 128 128
LR. schedule cosine decay cosine decay cosine decay cosine decay
LR. warmup ratio 0.03 0.03 0.03 0.03
Optimizer AdamW AdamW AdamW AdamW
Trainable Projector Full Full Full
modules Model Model Model
Table 9: 4-Stage training setup. We report performance after Stage-2 and Stage-3 of training in Tab. 6. Please note, both Stage-2 and Stage-3 are visual instruction tuning stages.
@@ 1-2B Model Comparison
Table 10: Breakdown of prefilling latencies for recent methods. The models are grouped based on total number of visual tokens. For models that were difficult to export or unavailable, we mark them as ’-’ in the table. “Vic.” refers to Vicuna [98], “Qw.2” refers to Qwen2 [86] and “Qw.” refers to Qwen [3]. “L-2” refers to LLaMA-2. “L-3” refers to LLaMA-3. “ML.” refers to MobileLLaMA [18, 19]. “DS.” refers to DeepSeek LLM [21]. For input resolution and visual tokens, we report the highest supported resolution by the respective models as some models like LLaVA-OneVision [45] and MM1 [66] use dynamic input resolution. FastVLM models using dynamic resolution employs a simple 2$\times$2 grid, with tile size set to 1024. For VLMs that use multiple vision encoders, the size of each encoder is listed independently, for TTFT, the latency from each encoder is summed up.
@@ C Additional Results
0.5B Model Comparison
R1 LLaVAOV [45] ViT-SO400M Qw.2 4.5+3.2 1152 7290 430 14124 61.4 - - 70.0 46.3
R2 FastVLM (Ours) FastViTHD Qw.2 15+12.5 1024 256 125 166 63.4 54.9 62.9 70.4 35.8
R3 FastVLM (Ours) FastViTHD Qw.2 15+23.1 1024 256 125 166 71.4 55.8 65.8 79.1 43.3
R4 FastVLM (Ours) FastViTHD Qw.2 15+12.5 2048 1280 125 918 68.8 59.0 65.4 82.1 49.3
R5 FastVLM (Ours) FastViTHD Qw.2 15+23.1 2048 1280 125 918 64.2 55.7 65.5 85.3 53.9
1-2B Model Comparison
R6 MM1 [66] ViT-H - 3000+1.5 1344 720 632 - 61.8 56.6 68.2 68.4 38.5
R7 FastVLM (Ours) FastViTHD Qw.2 15+12.5 1024 256 125 233 69.6 62.9 69.0 75.6 41.7
R8 FastVLM (Ours) FastViTHD Qw.2 15+23.1 1024 256 125 233 69.5 61.2 71.2 84.2 49.6
R9 FastVLM (Ours) FastViTHD Qw.2 15+12.5 2048 1280 125 1263 76.4 63.2 71.5 87.6 60.0
R10 FastVLM (Ours) FastViTHD Qw.2 15+12.5 2048 1280 125 1263 71.2 62.4 72.9 90.2 62.5
7B Model Comparison
R11 MM1 [66] ViT-H - 3000+1.5 1344 720 632 - 72.6 62.6 72.8 76.8 45.5
R12 LLaVA-NeXT ViT-L/14 L-3 - 672 2880 304 20347 69.5 49.0 64.6 72.6 -
R13 Cambrian-1 [78] ViT-L/14 L-3 2.5+7 336 576 304 5085 73.3 62.4 71.7 77.8 -
@@ 1-2B Model Comparison
R14 FastVLM (Ours) FastViTHD Vic. 0.5+0.6 768 144 125 387 17.1 30.0 62.9 32.9 28.7
R15 FastVLM (Ours) FastViTHD Vic. 0.5+1.1 768 144 125 387 59.1 38.4 67.5 57.3 29.7
R16 FastVLM (Ours) FastViTHD Vic. 15+1.1 768 144 125 387 65.4 45.3 69.4 65.5 32.0
R17 FastVLM (Ours) FastViTHD Qw.2 15+1.1 768 144 125 446 69.3 45.9 69.5 66.9 34.3
R18 FastVLM (Ours) FastViTHD Qw.2 15+11.9 768 144 125 446 74.2 59.0 72.8 72.0 44.3
R19 FastVLM (Ours) FastViTHD Vic. 0.5+0.6 1024 256 125 577 19.2 29.3 64.4 35.6 28.9
R20 FastVLM (Ours) FastViTHD Vic. 0.5+1.1 1024 256 125 577 61.0 38.3 67.4 62.8 32.0
R21 FastVLM (Ours) FastViTHD Vic. 15+1.1 1024 256 125 577 66.9 47.1 70.6 72.4 34.7
R22 FastVLM (Ours) FastViTHD Qw.2 15+1.1 1024 256 125 641 71.0 49.7 72.1 73.3 37.5
R23 FastVLM (Ours) FastViTHD Qw.2 15+6.5 1024 256 125 641 76.6 52.9 73.1 78.7 44.2
R24 FastVLM (Ours) FastViTHD Qw.2 15+11.9 1024 256 125 641 77.0 63.3 74.8 78.9 49.7
R25 FastVLM (Ours) FastViTHD Qw.2 15+12.5 1024 256 125 641 77.5 65.7 73.4 82.7 51.2
R26 FastVLM (Ours) FastViTHD Qw.2 15+23.1 1024 256 125 641 78.1 67.3 76.1 88.0 55.1
R27 FastVLM (Ours) FastViTHD Qw.2 15+12.5 2048 1280 125 3721 82.4 67.3 76.6 92.3 68.3
R28 FastVLM (Ours) FastViTHD Qw.2 15+23.1 2048 1280 125 3721 74.6 65.7 77.1 94.5 71.1
@@
Table 11: Comparison with recent methods on text-rich benchmarks. The models are grouped based on total number of visual tokens. “-” indicates that performance was not reported in the respective paper. For the dataset column, “-” indicates that the dataset size for pretraining (“PT”) or instruction tuning (“IT”) is not explicitly mentioned in the respective paper. For methods that have more than 2 stages of training, we report the total samples used for all the pretraining stages as part of “PT”. “TTFT” means time to first token (the sum of the vision encoder latency and the LLM prefilling time), we report latency only for models that are publicly available and in a format favorable to MLX [31] “Vic.” refers to Vicuna [98], “Qw.2” refers to Qwen2 [86]. “L-3” refers to LLaMA-3. * - For input resolution and visual tokens, we report the highest supported resolution by the respective models that use dynamic input resolution. †- performance numbers reported from [78]. For VLMs that use multiple vision encoders, the size of each encoder is listed independently, for TTFT, the latency from each encoder is summed up.
C.1 Dynamic Resolution (AnyRes) Results
From Fig. 6, it is evident that VLMs prefer visual encoding with fewer semantic breaks. Variants with more tiles typically underperform compared to those with fewer tiles and a static resolution. To scale up input resolution, we train variants of FastVLM with support for dynamic input resolution, where we use a tile size of 1024$\times$1024 and use a simple 2$\times$2 grid. This enables the model to process a peak input resolution of 2048$\times$2048 using only 4 tiles, unlike models like InternVL2 [16] which uses roughly 36 tiles to process images of resolution 2688$\times$2688. We report performance of FastVLM (R4, R5, R9, R10, R27, R28) with dynamic resolution support on text-rich benchmarks in Tab. 11.
C.2 CVBench and MathVista Results
Results in Tab. 12 show that, in comparison to Cambrian-1 [78], FastVLM is significantly better on MathVista [60] and competitive on CVBench [78], even though we have a single backbone and significantly fewer tokens. Results on both CVBench and MathVista benchmarks further improve as we scale the SFT dataset by including LLaVA-OneVision [45].
Method LLM Data (M) Input #Visual Latency CVBench CVBench Math
Decoder (PT+IT) Res. Tokens Enc.(ms) 2D 3D Vista
Cambrian-1 LLama3-8B 2.5+7 Mult. 576 3861.4 72.3 72.0 49.0
FastVLM Qwen2-7B 15+6.5 1024 256 116.3 71.8 69.6 57.6
FastVLM Qwen2-7B 15+11.9 768 144 54.8 75.9 79.3 64.6
FastVLM Qwen2-7B 15+11.9 1024 256 116.3 76.7 80.9 64.8
Table 12: Evaluation on CVBench and MathVista. “6.5M” instruction tuning dataset is from Cambrian-1. “11.9” instruction tuning dataset is concatenation of Cambrian-1 and LLaVA-OneVision datasets.